OpenStudy (yanasidlinskiy):

Help? Thanks! Please do not reply to this question. @johnweldon1993

OpenStudy (yanasidlinskiy):

\[\frac{ 21 }{ 35 } \div \frac{ 3 }{ 13 }\]

OpenStudy (yanasidlinskiy):

My answer was, 91/35. I don't know where I went wrong.

OpenStudy (johnweldon1993):

so wait....am I not to respond here? lol

OpenStudy (yanasidlinskiy):

Yes, respond here. I wanna knowxD

OpenStudy (johnweldon1993):

Lol I was like..."dont respond here" but then HOW!?!?!!? XD okay...you know that step 1 in dividing fractions...is to flip the second fraction and turn this to multiplication so \[\large \frac{21}{35} \times \frac{13}{3}\]

OpenStudy (yanasidlinskiy):

Lol, that was for other users....But yes. I know that. I flipped the second equation and did this: